A company’s capital consists of 100 000 ordinary shares issued at $2 and paid up to $1 per share. On 1 September, a first call of 50c was made on the ordinary shares. By 30 September, the call money received amounted to $45 000. No further payments were received, and on 31 October, the shares on which calls were outstanding were forfeited. On 15 November, the forfeited shares were reissued as paid to $1.50 for a payment of $1 per share. The appropriate cash amount from the reissue was received on 19 November. Costs of reissue amounted to $1 800. The company’s constitution provided for any surplus on resale, after satisfaction of unpaid calls, accrued interest and costs, to be returned to the shareholders whose shares were forfeited. The entry to record the reissue of forfeited shares is: A. Cash Dr 10 000 Forfeited shares Dr 5 000 Share capital — ordinary Cr 15 000 B. Cash Dr 5 000 Forfeited shares Dr 5 000 Share capital — ordinary Cr 10 000 C. Cash Dr 10 000 Share capital — ordinary Cr 10 000 D. Share capital Dr 15 000 Forfeited shares Cr 15 000

which one is true?
Answer #1

Answer - ( A )

Cash Dr 10,000 Forfeited shares Dr 5000 Share capital — ordinary Cr 15,000

EXPLANATION -

Entry for re-issue of shares

No

Journal

Debit

Credit

Working notes

1

Cash account

     10,000

[Actual amount received]
=10,000*1

Forfeited shares account

5,000

To Share capital

     15,000

[10,000*1.5]

Total money that should have been received on shares = 50c * 100,000shares = $50,000

Amount of money received = $45,000

Amount that is not received = $5,000

Number of shares on which the amount is not paid = $5,000/50c = 10,000 shares

If u r satisfied plz give a like

In case of any doubts plz let me know tanq